# Towards a Robust Transport Network With Self-adaptive Network Digital Twin

### v1

@@ 3.5 Network application
The network application is organized to use the predictions provided to the VTwin to perform adjustments in the PTwin considering applications that possibly will change the state of the PTwin, such as routing optimization or flow admission control or related to the management of the network, such as SLA monitoring for what-if analysis [6] or emergency preparedness [45]. To achieve this, various submodules can be employed within these applications, including, but not limited to, a recommendation system to propose different configurations in the PTwin, a policy evaluation module to evaluate the configurations suggested by the recommendation system, and a network management submodule designed to provide application-level metrics from the PTwin.
@@
In scenarios where the network application aims to optimize certain states of the PTwin, this module incorporates auxiliary tools, such as a recommendation system, that assist the network operator by leveraging outputs from the VTwin. This recommendation system can generate new policies that can be potentially applied to the PTwin. For example, in routing optimization, the recommendation system, which can be based on a genetic algorithm [27] or a reinforcement learning model [4], may propose new routing policies based on total per-flow delay metrics provided by the VTwin, aimed at optimizing the QoS metrics on the transport network. Once suggested, this recommendation is sent to the policy evaluation submodule, to be evaluated by a network operator, if a man-in-the-loop NDT is being used [2].
@@
In a man-in-the-loop architecture, a policy evaluation submodule is essential to prevent network modifications without human oversight, an important safeguard in production environments, particularly in architectures built on 5th generation (5G) networks [10]. This is especially critical given the interdependence of network components, such as the reliance of the radio access network on the transport network. The role of this submodule is to consolidate all relevant information and assess the potential impacts on the network, enabling informed decision-making by the operator. Once a policy is evaluated, the final decision on its implementation rests with the operator. If approved, the corresponding control data is transmitted to the SDN controller via the northbound API and subsequently applied to the PTwin, thus completing the NDT loop.
@@
When the application serves as a network management tool for what-if analysis, the task becomes conceptually simpler, as it does not require direct interaction with the PTwin to implement changes. Instead, its main objective is to provide the network operator with insight into the possible outcomes of specific scenarios or conditions [45], which can be useful in network planning tasks. This work explores a particular use case in the Experimental Results section, focusing exclusively on the network management submodule to provide SLA violations in the network [46, 47]. The selected application assesses compliance with flow SLAs, specifically detecting violations related to the packet delay budget (PDB) assigned to each packet, for example, if the total per-flow delay exceeds the PDB assigned for this flow. So, for a given set of flows that traverse the transport network, this application can estimate the number of SLA violations, relying on the total per-flow delay predictions provided by the VTwin.
@@ 4 Experiments
The PTwin framework integrates both emulation and simulation components to evaluate the proposed architecture. Using Mininet with open network operating system (ONOS) controller, the emulated component focuses on topology organization, while the simulated component handles traffic generation. For emulated network topologies, we use three transport networks: the PASSION project topology [48], the 5G-Crosshaul topology [48], and the Germany-50 topology [49], comprising networks with 128, 51, and 50 nodes, respectively, and each with 224, 88, and 63 links. Moreover, we employ the Distributed Internet Traffic Generator (D-ITG) using a client-server model over the Transmission Control Protocol (TCP) and User Datagram Protocol (UDP) protocols to generate realistic traffic patterns. To do so, we used a computer equipped with Intel® Core i7-10700F CPU @ 2.90 GHz, 128 GB of RAM, and a storage capacity of 9 TB. This setup enables the incorporation of various traffic distributions within an emulation environment that is orchestrated using a suite of tools, summarized in Table 3, to ensure accurate and reproducible testing conditions.
@@
We consider different distributions of packet size (in bytes) and traffic rate (in $\mathrm{M}\mathrm{b}\mathrm{i}\mathrm{t}\mathrm{s}\mathrm{/}\mathrm{s}$) distributions, without loss of generality, used to model internet traffic [50, 51], such as Poisson, Exponential, Uniform, and Deterministic, with the presence of congestion. The parameters for each are defined in Table 4, considering the minimum and maximum values $\mathcal{U}(512,1024)$ for the uniform distribution, the mean for the exponential $\mathcal{E}(1024)$ and $\mathcal{P}(2048)$ for a Poisson distribution, and a deterministic pattern with traffic congestion defined by a constant value $K=512$ for a constant packet size in traffic generation. For the packet rate, we considered the same distribution as defined for the packet size.
@@
Traffic pattern
Protocol
Packet size
@@
The datasets created from this traffic generation for a possible retraining process are presented in Table 5, characterized by the four distribution patterns of packet size and traffic rate, the type of topology, and the number of training examples. The training examples in this case refer to the number of samples that are to be used as input-output in the model retraining.
@@
Traffic pattern
Topology
Training
examples
1 Exponential 5G-Crosshaul 121 400121\,400121 400
Germany 129 600129\,600129 600
PASSION 129 600129\,600129 600
2 Poisson 5G-Crosshaul 234 501234\,501234 501
Germany 257 401257\,401257 401
PASSION 243 001243\,001243 001
3 Uniform 5G-Crosshaul 347 002347\,002347 002
Germany 378 802378\,802378 802
PASSION 355 002355\,002355 002
4
Deterministic
5G-Crosshaul 447 803447\,803447 803
Germany 440 003440\,003440 003
PASSION 478 203478\,203478 203
@@ 3 Uniform 5G-Crosshaul 347 002347\,002347 002
Considering these emulated environments into account, we propose an experiment using topologies that experience a sudden concept drift in traffic characteristics. We introduced an artificial drift at specific time intervals by altering the packet size distribution. As a result, the traffic follows the sequence of packet size and traffic rate patterns described below:
Start $\rightarrow \mathcal{E}$rightarrow $\mathcal{P}$rightarrow $\mathcal{U}$rightarrow $K$rightarrow End
@@
For each topology, we analyzed different time intervals of network traffic. Specifically, the total times in seconds for 5G-Crosshaul, Germany, and PASSION are 25 10325\,10325 103 $\mathrm{s}$, 24 60324\,60324 603 $\mathrm{s}$, and 26 70326\,70326 703 $\mathrm{s}$, respectively. Moreover, the experiment was carried out considering that we initially trained the VTwin model using traffic flows characterized by an exponential packet size distribution. The first abrupt drift is introduced by switching the packet size distribution to a Poisson model. This changing distribution behavior persists until the final phase, where the traffic pattern transitions to a constant packet size distribution, accompanied by the presence of traffic congestion.
Finally, we considered as a metric the NMSE to evaluate performance throughout these concept drifts, defined as:
$\text{NMSE}=10\log_{10}\left\{\dfrac{\sum\limits_{t=1}^{T}\left(y_{t}-\hat{y}_{t}\right)^{2}}{\sum\limits_{t=1}^{T}y_{t}^{2}}\right\},$ (7)
where $\hat{y}_{t}$ and $y_{t}$ represent the $t$-th estimated total per-flow delay and the corresponding ground truth total per-flow delay, respectively, across a total of $T$ examples in the test dataset. To compute the average, we used a window size of 100 consecutive examples over time. While other window sizes could be adopted, using 100 samples strikes a balance: it yields a sufficiently narrow confidence interval for the mean and enhances sensitivity in terms of NMSE to short-term variations, particularly important in the presence of abrupt concept drift.
@@
Using this environment, we present an experiment that considers a distribution-based detector aimed at detecting input changes. We configured the concept drift detector using the Kolmogorov-Smirnov windowing (KSWIN) method [52], as implemented by the river library [53]. To work correctly, this method requires three hyperparameters: the significance level, which is a sensitive parameter that controls the probability of false positives, represented by a value of 0.001, a window size, and a buffer for the most recent samples. The sizes of these components vary across different topologies and were determined empirically.
For the VTwin training, we used TensorFlow 2.14.0 and maintained consistency in all training parameters in all experiments. Specifically, we used a learning rate of 0.0010.0010.001, trained for 50 epochs, applied Z-score normalization for feature scaling, adopted the Adam optimizer, and used Mean Absolute Percentage Error (MAPE) as the loss function. Moreover, as the model hyperparameters, we used: $T=12$ and $||\mathbf{h}_{f}||=||\mathbf{h}_{l}||=16$. This uniform configuration was chosen to ensure that any observed differences in model performance could be attributed solely to variations in the input data.
Experimental Results
We have the experiments from the KSWIN results to evaluate the proposed synchronization elements, considering the three emulated topologies and the average traffic rate in $\mathrm{M}\mathrm{b}\mathrm{i}\mathrm{t}\mathrm{s}\mathrm{/}\mathrm{s}$ as the feature to be tracked. Figure 6 illustrates the operation of the KSWIN detector and the location of each detected drift. For each network topology, we selected a specific window size (in seconds): 6800 for 5G-Crosshaul, 7000 for Germany, and 6800 for the PASSION topology. Although alternative window sizes could be explored, determining the optimal size to minimize false positives falls outside the scope of this work.
Figure 6: Drift detection pattern considering 5G-Crosshaul, Germany and PASSION topologies and four type of traffic patterns.
In the evaluated environments, smaller window sizes generally result in a higher likelihood of false positives, as illustrated in Figure 7, which analyzes various window sizes in all topologies. The curves in this figure exhibit a consistent pattern and converge in the number of detected concept drifts as the window size increases. This behavior can be attributed to the timing of the concept drift events. For example, the first concept drift, shown in Figure 6, occurs approximately in the second 7000, across all topologies, coincident with the point at which the curves in Figure 7 end converging.
Figure 7: Relationship between the window size and the number of concept drift detection on the three topologies.
In this experiment, we can observe that concept drift was effectively detected in all existing deviations due to changes in the traffic pattern. However, there were different delays, considering the time when the concept drift occurred and when the detection was made. Considering the effects of these delayed drift detections, we can see in Figure 8 the VTwin performance over time considering the proposed synchronization elements in comparison with the same VTwin without them in terms of NMSE.
Figure 8: Validating proposed architecture on 5G-Crosshaul topology. The blue line represents the VTwin performance with the proposed synchronization elements, while the gray dashed line represents the performance without it.
Before the first concept drift is detected, the VTwin operates in a typical manner, as seen in the related work: a GNN model already trained to generalize to unseen datasets and, based on its predictions, performs the required analysis within the proposed application. However, after the first concept drift occurs, the scenario changes due to traffic variability, leading to model degradation, evidenced by some sample windows exhibiting NMSE values near 0 $\mathrm{d}\mathrm{B}$. Despite this decline, the VTwin model remains stable, even without retraining, as the underlying data distributions, such as Poisson and Exponential, are still related. This behavior changes after the second concept drift, as the input data distribution changes to a packet size with a distinct pattern, requiring retraining, which is evident in the accentuated NMSE difference between the model with and without retraining. In the last concept drift, the VTwin becomes obsolete without a retraining process as the input data distribution shifts to a UDP connection characterized by constant packet size and the absence of congestion control. At this point, retraining becomes extremely necessary, as demonstrated by the pronounced performance gap between the models with and without retraining.
The results for the other two topologies aim to validate the purpose synchronization method considering topologies with different characteristics, for instance, in the PASSION topology that has 128 nodes and three times more links than the 5G-crosshaul. In this way, Figure 9 also shows consistent results in both topologies as those observed in the 5G-Crosshaul topology, specifically in the last drift detected by the concept drift detector, highlighting the same observations related to mandatory retraining due to poor VTwin prediction performance.
Figure 9: Evaluation performance of proposed architecture considering Germany and PASSION topologies.
@@
Table 6: NMSE performance in $\mathrm{d}\mathrm{B}$ before and after concept drift, considering the NDT operation with and without the proposed synchronization module implementation
Topology Drift ID
Pre-drift avg.
NMSE w/o sync.
Post-drift avg.
NMSE w/o sync.
Pre-drift avg.
NMSE w/ sync.
Post-drift avg.
NMSE w/ sync.
5G-Crosshaul 1 20.72-20.72- 20.72 10.84-10.84- 10.84 20.72-20.72- 20.72 $\mathbf{-14.47}$
2 10.84-10.84- 10.84 5.42-5.42- 5.42 14.47-14.47- 14.47 $\mathbf{-11.22}$
3 5.42-5.42- 5.42 7.787.787.78 11.22-11.22- 11.22 $\mathbf{1.43}$
Germany 1 19.98-19.98- 19.98 12.11-12.11- 12.11 19.98-19.98- 19.98 $\mathbf{-20.10}$
2 12.11-12.11- 12.11 1.50-1.50- 1.50 20.10-20.10- 20.10 $\mathbf{-14.85}$
3 1.50-1.50- 1.50 25.1025.1025.10 14.85-14.85- 14.85 $\mathbf{-5.73}$
PASSION 1 18.44-18.44- 18.44 6.62-6.62- 6.62 18.44-18.44- 18.44 $\mathbf{-19.17}$
2 6.62-6.62- 6.62 0.470.470.47 19.17-19.17- 19.17 $\mathbf{-8.04}$
3 0.470.470.47 38.7038.7038.70 8.04-8.04- 8.04 $\mathbf{5.83}$

### v2

@@ 3.5 Network application
The network application is organized to use the predictions provided to the VTwin to perform adjustments in the PTwin, considering applications that may change the state of the PTwin, such as routing optimization or flow admission control, as well as those related to the management of the network, such as SLA monitoring for what-if analysis [Wang2022] or emergency preparedness [TR28915]. To achieve this, various submodules can be employed within these applications, including, but not limited to, a recommendation system to propose different configurations in the PTwin, a policy evaluation module to evaluate the configurations suggested by the recommendation system, and a network management submodule designed to provide application-level metrics from the PTwin.
@@
In scenarios where the network application aims to optimize certain states of the PTwin, this module incorporates auxiliary tools, such as a recommendation system, that assist the network operator by leveraging outputs from the VTwin. This recommendation system can generate new policies that can potentially be applied to the PTwin. For example, in routing optimization, the recommendation system, which can be based on a genetic algorithm [Zalat2024] or a reinforcement learning model [abenathar2025], may propose new routing policies based on total per-flow delay metrics provided by the VTwin, aimed at optimizing the QoS metrics on the transport network. Once suggested, this recommendation is sent to the policy evaluation submodule to be evaluated by a network operator if a human-in-the-loop NDT is being used [ohlen2022, abenathar2025].
@@
In a man-in-the-loop architecture, a policy evaluation submodule is essential to prevent network modifications without human oversight, which is an important safeguard in production environments, particularly in architectures built on 5th generation (5G) networks [messaoudi2023]. This is especially critical given the interdependence of network components, such as the reliance of the RAN on the transport network. The role of this submodule is to consolidate all relevant information and assess the potential impacts on the network, enabling informed decision-making by the operator. Once a policy is evaluated, the final decision on its implementation rests with the operator. If approved, the corresponding control data is transmitted to the SDN controller via the northbound API and subsequently applied to the PTwin, thus completing the NDT loop.
@@
When the application serves as a network management tool for what-if analysis, the task becomes conceptually simpler, as it does not require direct interaction with the PTwin to implement changes. Instead, its main objective is to provide the network operator with insight into the possible outcomes of specific scenarios or conditions [TR28915], which can be useful in network planning tasks. This work explores a particular use case in the Experimental Results section, focusing exclusively on the network management submodule to provide SLA violations in the network [Sami2022, sommers2010]. The selected application assesses compliance with flow SLAs, specifically detecting violations related to the packet delay budget (PDB) assigned to each packet. For example, if the total per-flow delay exceeds the PDB assigned for this flow. So, for a given set of flows that traverse the transport network, this application can estimate the number of SLA violations, relying on the total per-flow delay predictions provided by the VTwin.
@@ 4 Experiments
The PTwin integrates both emulation and simulation components used as part of the experiments to evaluate the proposed architecture. Using Mininet with the Open Network Operating System (ONOS) controller, the emulated component focuses on topology organization, while the simulated component handles traffic generation. For emulated network topologies, we use four transport networks: the 5G-Crosshaul topology [Morais2023], the PASSION project topology [Morais2023], the Germany topology [kdn_datasets], and the Synthetic-700 topology comprising networks with 128, 51, 50, and 700 nodes, respectively, each with 63, 88, 224, and 1083 links. The details about the characteristics of link capacity and link propagation delay are also described in Figures 9(a) and  9(b), respectively.
(a) Histogram of link capacities.
(b) Histogram of link propagation delay.
Figure 9: Link characteristics of the topologies used in our experiments.
We employ the Distributed Internet Traffic Generator (D-ITG) using a client-server model over the Transmission Control Protocol (TCP) and User Datagram Protocol (UDP) protocols to generate realistic traffic patterns. To do so, we used a computer equipped with an Intel® Core i7-10700F CPU @ 2.90 GHz, an NVIDIA® GeForce RTX 3060, 128 GB of RAM, and a storage capacity of 9 TB. This setup enables the incorporation of various traffic distributions within an emulation environment that is orchestrated using a suite of tools summarized in Table 4 to ensure accurate and reproducible testing conditions.
@@
We consider different distributions of packet size (in bytes) and traffic rate (in $\mathrm{M}\mathrm{b}\mathrm{i}\mathrm{t}\mathrm{s}\mathrm{/}\mathrm{s}$), without loss of generality, used to model internet traffic [Jurkiewicz2021, Song2005], such as Poisson, Exponential, Uniform, and Deterministic, in the presence of congestion. The parameters for each are defined in Table 5, considering the minimum and maximum values $\mathcal{U}(512,1024)$ for the uniform distribution, the mean for the exponential $\mathcal{E}(1024)$ and $\mathcal{P}(2048)$ for a Poisson distribution, as well as a deterministic pattern with traffic congestion defined by a constant value $K=512$ for a constant packet size in traffic generation. For the packet rate, we considered the same distribution as defined for the packet size. This simplification regarding the packet rate distribution was made to maintain a controlled environment and to gain a clearer understanding of the strengths and limitations of the proposed architecture.
@@
Traffic pattern Protocol Packet size (bytes)
@@
The datasets created from this traffic generation for a possible retraining process are presented in Table 6, characterized by the four distribution patterns of packet size and traffic rate, the type of topology, and the number of flows (training examples). The training examples in this case refer to the number of samples that are to be used as input-output pairs in the model retraining.
@@
# Traffic pattern Topology Number of flows ($T$)
1 Exponential 5G-Crosshaul 121 400121\,400
Germany 129 600129\,600
PASSION 129 600129\,600
Synthetic-700 170 000170\,000
2 Poisson 5G-Crosshaul 113 100113\,100
Germany 127 800127\,800
PASSION 113 400113\,400
Synthetic-700 194 400194\,400
3 Uniform 5G-Crosshaul 112 500112\,500
Germany 121 400121\,400
PASSION 112 000112\,000
Synthetic-700 162 000162\,000
4 Deterministic 5G-Crosshaul 100 800100\,800
Germany 61 20061\,200
PASSION 123 200123\,200
Synthetic-700 186 100186\,100
@@ 4 Deterministic 5G-Crosshaul 100 800100\,800
Considering these emulated environments, we propose several experiments using topologies that experience a sudden concept drift in traffic characteristics. We introduced an artificial drift at specific time intervals by altering the packet size distribution. As a result, the traffic follows the sequence of packet size and traffic rate patterns described below:
Start $\rightarrow$ $\mathcal{E}$ $\rightarrow$ $\mathcal{P}$ $\rightarrow$ $\mathcal{U}$ $\rightarrow$ $K$ $\rightarrow$ End
@@
For each topology, we analyzed different time intervals of network traffic. Specifically, the total times in seconds for 5G-Crosshaul, Germany, PASSION, and Synthetic-700 are 25 10025\,100 $\mathrm{s}$, 24 60024\,600 $\mathrm{s}$, 26 70026\,700 $\mathrm{s}$, and 39 700 $\mathrm{s}$, respectively. Moreover, the experiment was carried out with the understanding that we initially trained the VTwin model using traffic flows characterized by an exponential packet size distribution. The first abrupt drift is introduced by switching the packet size distribution to a Poisson model. This changing distribution behavior persists until the final phase, where the traffic pattern transitions to a constant packet size distribution, accompanied by the presence of traffic congestion. Finally, we considered the NMSE as a metric to evaluate the VTwin performance with and without NDT synchronization throughout these concept drifts, defined as:
$\text{NMSE}^{(q)}=10\log_{10}\left\{\dfrac{\sum\limits_{t=1}^{T}\left(y_{t}^{(q)}-\hat{y}_{t}^{(q)}\right)^{2}}{\sum\limits_{t=1}^{T}\left(y_{t}^{(q)}\right)^{2}}\right\},$ (8)
where $\hat{y}_{t}^{(q)}$ and $y_{t}^{(q)}$ denote, respectively, the $t$-th predicted and ground-truth QoS metric among $T$ test samples. The average values were computed using a sliding window of 100 consecutive samples over time. Although other window sizes could be used, a 100-sample window offers a suitable trade-off: it provides a sufficiently narrow confidence interval for the mean while maintaining high NMSE sensitivity to short-term variations, an aspect particularly relevant during abrupt concept drift. We apply this evaluation across 10 independent realizations, aimed at obtaining an average perspective across the NDT operation in each network topology.
@@
Using this environment, our experiments consider a distribution-based detector aimed at detecting changes in the input. We configured the concept drift detector using the Kolmogorov-Smirnov windowing (KSWIN) method [raab2020], as implemented by the river library [montiel2021river]. To work correctly, this method requires three hyperparameters: the significance level, which is a sensitive parameter that controls the probability of false positives and is represented by a value of 0.001; a window size; and a buffer for the most recent samples. The sizes of these components vary across different topologies and are determined empirically.
@@
Figure 10: Experiment organization to evaluate the VTwin performance with and without synchronization in a scenario with concept drift inducted by different traffic pattern.
Concept drift detector configuration
We have the experiments from the KSWIN results to evaluate the proposed synchronization elements, considering the four emulated topologies and the average traffic rate in $\mathrm{M}\mathrm{b}\mathrm{i}\mathrm{t}\mathrm{s}\mathrm{/}\mathrm{s}$ as the features to be tracked. Figure 11 illustrates the operation of the KSWIN detector and the location of each detected drift. For each network topology, we selected a specific window size (in seconds): 6 8006\,800 for 5G-Crosshaul, 7 0007\,000 for Germany, 6 8006\,800 for PASSION, and 10 00010\,000 for the Synthetic-700 topology. Although alternative window sizes could be explored, determining the optimal size to minimize false positives falls outside the scope of this work.
Figure 11: Drift detection pattern considering 5G-Crosshaul, Germany, PASSION, and Synthetic-700 topologies and four type of traffic patterns. Across the time series, each color in a given interval represents a different distribution pattern. The “x” points represent the detected concept drift.
In the evaluated environments, smaller window sizes generally result in a higher likelihood of false positives, as illustrated in Figure 12, which analyzes various window sizes across all topologies. The curves in this figure exhibit a consistent pattern and converge in the number of detected concept drifts as the window size increases.
Figure 12: Relationship between the window size and the number of concept drift detection on the four topologies.
Delay prediction performance
In the first experiment, we observed that concept drift was successfully detected in all instances of deviation caused by changes in the traffic pattern. However, varying detection delays were identified between the moment when the drift occurred and the moment it was detected. The resulting performance is presented in Figure 13, which shows the VTwin performance over time in the 5G-Crosshaul and Germany topology in terms of NMSE. The figure compares our proposed synchronization elements against a baseline VTwin without them, where the baseline scenario can be interpreted as one in which every concept drift results in a false negative.
Figure 13: Validation of the proposed architecture in the 5G-Crosshaul and Germany topologies. The blue line represents the average per-flow delay prediction performance of 10 retraining realizations (in terms of NMSE) with the proposed synchronization elements. The gray dashed line shows the performance without them, and the light blue shaded area indicates the standard deviation of NMSE across all retraining realizations. Finally, the red-shaded area represents the region in which the retraining process occurred across realizations.
Before the first concept drift is detected, the VTwin operates in a typical manner, as seen in the related work: a GNN model that is already trained to generalize to unseen datasets and, based on its predictions, performs the required analysis within the proposed application. However, after the first concept drift occurs, the scenario changes due to traffic variability, leading to model degradation, evidenced by some sample windows exhibiting NMSE values near 0 $\mathrm{d}\mathrm{B}$. Despite this decline, the VTwin model remains stable, even without retraining, as the underlying data distributions, such as Poisson and Exponential, are still related. This behavior changes after the second concept drift, as the input data distribution changes to a packet size with a distinct pattern, requiring retraining, which is evident in the accentuated NMSE difference between the models with and without retraining. In the last concept drift, the VTwin becomes obsolete without a retraining process, as the input data distribution shifts to a UDP connection characterized by a constant packet size and the absence of congestion control. At this point, retraining becomes extremely necessary, as demonstrated by the pronounced performance gap between the models with and without retraining.
The results obtained for the remaining two topologies are intended to further validate the effectiveness of the proposed synchronization method across networks with diverse characteristics in terms of scalability. Specifically, the PASSION topology comprises 128 nodes and approximately three times more links than the 5G-Crosshaul topology, while the Synthetic-700 topology contains roughly thirteen times more nodes than the Germany topology. Despite these substantial differences in scale and connectivity, Figure 14 shows consistent behavior across both topologies, comparable to that observed in the 5G-Crosshaul and Germany scenarios. In particular, during the final concept drift detected by the drift detector, the results reinforce the same conclusions regarding the necessity of mandatory retraining due to degraded VTwin prediction performance.
Figure 14: Comparison of the per-flow delay prediction performance using the proposed architecture considering PASSION and Synthetic-700 topologies.
In terms of retraining time cost in minutes, across these 10 realizations, we obtained the following results depicted by Figure 15. In our environment, the training time analysis shows that three out of four retraining processes for the delay prediction task are completed in less than $35$ minutes. The exception is the Synthetic-700 topology, which is explained by the size of the topology. Considering the overall NDT operation time, the feasibility of the retraining duration depends on the specific application, and optimizing this time lies beyond the scope of this work. In our scenario, for instance, within the 5G-Crosshaul topology, 83.41 minutes are spent in retraining states out of a total of 418.8 minutes, corresponding to approximately 20% of the total operation time. For the Germany and PASSION topologies, the proportion of time devoted to retraining remains below 22%. Finally, for the Synthetic-700, the retraining time spent was below 32%, indicating a reasonable balance between adaptability and operational continuity.
Noteworthy, in certain concept drift events, the retraining time can be considerably shorter. For instance, in the PASSION topology, following the second detected drift, the average retraining duration was only $9.42$ minutes. This reduction is primarily due to the use of an early stopping callback, which prevents unnecessary training of the VTwin model once convergence is reached.
Figure 15: Average retraining time (in minutes) cost by the VTwin to synchronize with the PTwin in the delay prediction task. The value above each bar represents the average retraining time.
Finally, Table 7 summarizes the NMSE across all window examples, highlighting VTwin’s performance before and after each concept drift. For the NMSE after the concept drift, we assume the mean average NMSE for comparison with the approach without synchronization, since our evaluation considers 10 retraining process realizations. Hence, incorporating our NDT synchronization module results in a performance improvement of at least $64\%$ in predictions compared to configurations without it. Furthermore, given our empirical goal of achieving an average NMSE below 10-10 $\mathrm{d}\mathrm{B}$, the approach without NDT synchronization meets this target in only 2 of the 12 concept drifts across all evaluated topologies. In contrast, our proposed approach with NDT synchronization achieves this performance in 6 out of 12 cases following concept drift. Thus, from a general perspective, for all topologies and after all concept drifts, the performance of the predictions has improved, highlighting the importance of twin synchronization in production environments where concept drift is pervasive.
@@
Table 7: NMSE performance in predicting per-flow delay (in $\mathrm{d}\mathrm{B}$) before and after concept drift, considering the NDT operation both with and without the proposed synchronization module.
Topology Drift ID Pre-drift avg. NMSE w/o sync. Post-drift avg. NMSE w/o sync. Pre-drift mean avg. NMSE w/ sync. Post-drift mean avg. NMSE w/ sync.
5G-Crosshaul 1 20.69-20.69 14.36-14.36 20.69-20.69 $\mathbf{-18.87}$
2 14.36-14.36 5.11-5.11 18.87-18.87 $\mathbf{-10.10}$
3 5.11-5.11 $17.40$ 10.10-10.10 $\mathbf{2.08}$
Germany 1 18.88-18.88 14.10-14.10 18.88-18.88 $\mathbf{-19.01}$
2 14.10-14.10 1.79-1.79 19.01-19.01 $\mathbf{-13.82}$
3 1.79-1.79 $24.64$ 13.82-13.82 $\mathbf{-3.27}$
PASSION 1 19.95-19.95 5.41-5.41 19.95-19.95 $\mathbf{-19.01}$
2 5.41-5.41 1.14-1.14 19.01-19.01 $\mathbf{-8.06}$
3 1.14-1.14 $38.09$ 8.06-8.06 $\mathbf{7.19}$
Synthetic-700 1 3.76-3.76 1.44-1.44 3.76-3.76 $\mathbf{-13.77}$
2 1.44-1.44 0.93-0.93 13.77-13.77 $\mathbf{-6.42}$
3 0.93-0.93 $14.89$ 6.42-6.42 $\mathbf{-2.05}$
Jitter prediction performance
Using the same environment, we also evaluated the per-flow jitter prediction performance in scenarios with and without NDT synchronization. In this sense, Figures 16 and 17 depict the performance in terms of NMSE in the 5G-Crosshaul, Germany, PASSION, and Synthetic-700 topologies. From these results, it is possible to perceive similar trends with NDT synchronization. For example, after the third detected concept drift, the retraining process is extremely necessary.
Figure 16: Comparison of the jitter prediction performance using the proposed architecture considering 5G-Crosshaul and Germany topologies.
Figure 17: Comparison of the jitter prediction performance using the proposed architecture considering PASSION and Synthetic-700 topologies.
Furthermore, a similar pattern to that observed in the delay prediction training cost can be seen in the average retraining time after a concept drift event for the jitter prediction task, as shown in Figure 18. In this case, with the exception of experiments with the Synthetic-700 topology, all retraining times remained below $35$ minutes, resulting in a retraining time proportion below $20\%$ for 5G-Crosshaul, Germany, PASSION, and below $30\%$ for the Synthetic-700 topology.
Figure 18: Average retraining time (in minutes) cost by the VTwin to synchronize with the PTwin in the jitter prediction task. The value above each bar represents the average retraining time.
These results for the jitter prediction task reaffirm the important role of the proposed NDT synchronization architecture in maintaining satisfactory performance on QoS prediction. As shown in Table 8, when using our synchronization approach, we achieved an NMSE below $10~$\mathrm{d}\mathrm{B}$\$ in 8 out of 12 drift events, whereas the scenario without synchronization reached this level in only one event.
Table 8: NMSE performance in predicting per-flow jitter (in $\mathrm{d}\mathrm{B}$) before and after concept drift, considering the NDT operation both with and without the proposed synchronization module.
Topology Drift ID Pre-drift avg. NMSE w/o sync. Post-drift avg. NMSE w/o sync. Pre-drift mean avg. NMSE w/ sync. Post-drift mean avg. NMSE w/ sync.
5G-Crosshaul 1 19.78-19.78 5.84-5.84 19.78-19.78 $\mathbf{-13.66}$
2 5.84-5.84 7.60-7.60 13.66-13.66 $\mathbf{-8.67}$
3 7.60-7.60 1.44-1.44 8.67-8.67 $\mathbf{-3.83}$
Germany 1 18.30-18.30 5.23-5.23 18.30-18.30 $\mathbf{-12.01}$
2 5.23-5.23 7.63-7.63 12.01-12.01 $\mathbf{-11.43}$
3 7.63-7.63 $7.88$ 11.43-11.43 $\mathbf{-9.46}$
PASSION 1 23.10-23.10 4.17-4.17 23.10-23.10 $\mathbf{-15.78}$
2 4.17-4.17 10.32-10.32 15.78-15.78 $\mathbf{-15.68}$
3 10.32-10.32 $6.37$ 15.68-15.68 $\mathbf{-12.98}$
Synthetic-700 1 20.89-20.89 6.39-6.39 20.89-20.89 $\mathbf{-11.02}$
2 6.39-6.39 4.30-4.30 11.02-11.02 $\mathbf{-7.55}$
3 4.30-4.30 $11.23$ 7.55-7.55 $\mathbf{-11.51}$